OpenStudy (tparker22):

How many atoms are in 0.8288 grams of sodium, Na?

OpenStudy (holsteremission):

\[0.8288\text{ g}\times\frac{1\text{ mol}}{22.99\text{ g}}\times\frac{6.022\times10^{23}\text{ atoms}}{1\text{ mol}}=\cdots\text{ atoms}\]